# 3DGS-DET: Empower 3D Gaussian Splatting with Boundary Guidance and Box-Focused Sampling for 3D Object Detection

### v1

@@ 4.1 Experimental setup
Dataset: To thoroughly evaluate the performance of our proposed method in 3D detection tasks, we selected two representative datasets: ScanNet (Dai et al., 2017) and ARKitScene (Baruch et al., 2021). ScanNet is a large-scale indoor scene dataset containing over 1,500 real-world 3D scanned scenes, encompassing various complex indoor environments such as residential spaces, offices, and classrooms. The ARKitScene dataset is constructed from RGB-D image sequences, offering detailed geometric information and precise object annotations. For each scene, a maximum of 600 posed images are extracted. The category settings follow the standard 18 categories for ScanNet and 17 categories for ARKitScene.
Metrics: We use mAP@0.25 and mAP@0.5 as the primary evaluation metrics. Mean Average Precision (mAP) is calculated at different IoU thresholds, providing a comprehensive measure of the detection model’s performance across various categories.
Implementation Details: For training 3DGS, we follow Kerbl et al. (2023) to initialize the 3D coordinates of Gaussian blobs using Structure from Motion (SfM) points. The training hyperparameters are the same as those in Kerbl et al. (2023). We employ pretrained GroundedSAM (Ren et al., 2024) and the Suzuki-Abe algorithm (Suzuki et al., 1985) as the boundary detector in Boundary Guidance. The pretrained GroundingDINO (Liu et al., 2023) is used as the box detector in the Box-Focused Sampling strategy. For the detection tool, we utilize the FCAF3D (Rukhovich et al., 2022a) architecture implemented in MMDetection3D (Contributors, 2020). The training hyperparameters are the same as those in FCAF3D. In our ablation study, to ensure a fair comparison, all model versions are trained with the same hyperparameters, such as the same number of epochs, specifically 12 epochs. All the ablation experiments (Sec. 4.3) are conducted on ScanNet.
4.2 Main Results
Quantitative Results. For the ScanNet dataset, we present the mAP@0.25 and mAP@0.5 performances of various methods in Tab. 1 of the main paper and Tab. 6 of the Appendix, respectively. Note that some methods did not report mAP@0.5 in previous studies, resulting in blank entries for these methods in Tab. 6 of the Appendix.
In both Tab. 1 and Tab. 6 of the Appendix, the methods listed in the first block (Wang et al., 2018; He et al., 2017; Hou et al., 2019; Qi et al., 2019; Rukhovich et al., 2022a; Wang et al., 2022a; Tu et al., 2023; Shen et al., 2024a; Rukhovich et al., 2022b) are non-view-synthesis representation-based 3D detection methods. These methods utilize point clouds, RGB-D data, or multi-view images for 3D object detection. The second block consists of view-synthesis representation-based 3DOD methods, including NeRF-Det (Hu et al., 2023) and our proposed 3DGS-DET. NeRF-Det is the closest work to ours, leveraging Neural Radiance Fields (NeRF). Our approach variants are detailed as follows: ‘3DGS-DET (Our basic pipeline)’ represents the basic pipeline method established in Sec. 3.2. ‘3DGS-DET (Our basic pipeline+BG)’ incorporates the proposed Boundary Guidance as detailed in Sec. 3.3. ‘3DGS-DET (Our basic pipeline+BG+BS)’ is our full method, utilizing both Boundary Guidance and Box-Focused Sampling as described in Sec. 3.4. As illustrated in Tab. 1 and Tab. 6 of the Appendix, all versions of our methods significantly outperform NeRF-Det. Notably, our full method (’Our basic pipeline+BG+BS’) surpasses the state-of-the-art NeRF-based method, NeRF-Det, by +6.6 on mAP@0.25 and +8.1 on mAP@0.5, showcasing the superiority of our approach.
Table 2: Comparison of the ‘whole-scene’ performance on the ARKITScenes validation set. Our 3DGS-DET significantly outperforms NeRF-Det by 31.5 points. Note that we follow the setup described in the NeRF-Det (Xu et al., 2023) supplementary materials: ‘In our experiments, we utilize the subset of the dataset with low-resolution images’, considering it is the closest work to ours. Other methods that do not use the same setting are not listed in this table.
Methods cab fridg shlf stove bed sink wshr tolt bthtb
ImVoxelNet (Rukhovich et al., 2022b) 32.2 34.3 4.2 0.0 64.7 20.5 15.8 68.9 80.4
NeRF-Det (Xu et al., 2023) 36.1 40.7 4.9 0.0 69.3 24.4 17.3 75.1 84.6
3DGS-DET (Ours) 45.2 84.4 33.3 41.4 87.3 75.5 67.6 87.2 90.8
Methods oven dshwshr frplce stool chr tble TV sofa mAP@.25
ImVoxelNet (Rukhovich et al., 2022b) 9.9 4.1 10.2 0.4 5.2 11.6 3.1 35.6 23.6
NeRF-Det (Xu et al., 2023) 14.0 7.4 10.9 0.2 4.0 14.2 5.3 44.0 26.7
3DGS-DET (Ours) 74.3 6.0 56.4 26.3 70.3 60.6 0.7 81.8 58.2  (+31.5)
Regarding the ARKitScene dataset, considering NeRF-Det is the closest work to ours, we follow the same setup described in the NeRF-Det (Xu et al., 2023) supplementary materials: ‘In our experiments, we utilize the subset of the dataset with low-resolution images.’ Similarly, we adopt the same subset of the ARKitScenes dataset. Other methods that report performance on ARKitScene use the full dataset, so our 3DGS-DET is only compared with ImVoxelNet and NeRF-Det under the same conditions as described in NeRF-Det. The results in Tab. 2 demonstrate that 3DGS-DET performs better across most categories, achieving an mAP@0.25 of 58.2, which significantly outperforms NeRF-Det by +31.5, highlighting the superiority of our method.
Qualitative results. We provide a qualitative comparison with NeRF-Det in Fig. 4. As shown, our methods detect more objects in the scene with greater positional accuracy compared to NeRF-Det (Xu et al., 2023), demonstrating the superiority of our approach. More qualitative comparisons can be found in Fig. 6 and Fig. 7 in the Appendix.
@@ 4.2 Main Results
Figure 4: Qualitative comparison. Our methods identify more 3D objects in the scene with better positional precision, highlighting the advantages of our approach over NeRF-Det (Xu et al., 2023). In this figure, the scene is represented using mesh to clearly show the boxes.
4.3 Ablation Study
4.3.1 Analysis on the Effect of Proposed Designs
In this section, we demonstrate the effectiveness of our contributions by first presenting the performance of our proposed basic 3DGS detection pipeline and then incrementally incorporating our additional designs to analyze the resulting performance improvements.
Our Proposed Basic 3DGS Detection Pipeline. As shown in Tab. 1, ’3DGS-DET (Our basic pipeline)’ represents our proposed detection pipeline utilizing 3DGS, as described in Section 3.2. Benefiting from the advantages of 3DGS as an explicit scene representation, our basic pipeline surpasses NeRF-Det by 1 point (54.3 vs. 53.3), underscoring the significance of introducing 3DGS into 3DOD for the first time.
Table 3: Ablation study on guidance from different priors.
@@ 4.3.1 Analysis on the Effect of Proposed Designs
Table 4: Ablation study on different sampling methods.
@@
Boundary Guidance. ‘3DGS-DET (Our basic pipeline+BG)’ incorporates the proposed Boundary Guidance as detailed in Sec. 3.3. Introducing Boundary Guidance into the basic pipeline results in a significant improvement of 2.4 points (56.7 vs. 54.3), demonstrating the effectiveness of the proposed Boundary Guidance. To further explore the impact of Boundary Guidance on 3DGS representations, we present a visual comparison of the spatial distribution of trained Gaussian blobs in Fig. 8 in the Appendix. As we can see, Gaussian blobs trained with Boundary Guidance demonstrate clearer spatial distribution and more distinct differentiation between objects and the background. We also present rendered images from different views by 3DGS trained with Boundary Guidance in Fig. 9 and Fig. 10 in the Appendix. As can be observed, the category-specific boundaries are clearly rendered and show multi-view stability, indicating that the 3D representation has effectively embedded the priors from Boundary Guidance. All these results clearly verify the effectiveness of the proposed Boundary Guidance for 3D detection with 3DGS.
@@
Box-Focused Sampling. Furthermore, we introduce Box-Focused Sampling detailed in Sec. 3.4, represented by ‘3DGS-DET (Our basic pipeline+BG+BS)’ in Tab. 1. This addition leads to a further performance boost of 3.2 points (59.9 vs. 56.7), proving the effectiveness of Box-Focused Sampling. The visual comparison of sampled Gaussian blobs is shown in Fig. 11 in the Appendix. We can observe that the proposed Box-Focused Sampling significantly retains more object blobs and suppresses noisy background blobs.
[…]
Figure 5: Analysis of guidance from different priors: (a) Center Point Guidance, (b) Mask Guidance, and (c) Boundary Guidance. In (a) and (b), the spatial distribution of Gaussian blobs for objects like the chair, trash bin and sink is incomplete and ambiguous. Gaussian blobs trained with Boundary Guidance exhibit a clearer spatial distribution. The reason behind this phenomenon is that the center point provides only positional guidance, lacking richer information like shape or size. The mask highlights shape and size but hides the object’s surface, reducing texture and geometric information. Boundary Guidance offers positional cues and richer information, such as shape and size, while preserving texture and geometric details on the object’s surface, leading to the best performance.
4.3.2 Ablation Study on Guidance from Different Priors
@@
In this section, we analyze the impact of guidance from various priors. As described in Sec. 3.3, we utilize the object’s boundary as the guidance prior. Here, we perform an ablation study considering the object’s center point and mask as alternative priors. To obtain the center point, we detect the object’s bounding box using GroundingDINO (Liu et al., 2023) and compute its center coordinates. The mask is generated with GroundedSAM (Ren et al., 2024). Note that all priors are category-specific, with each class associated with a fixed color. These priors are overlaid on the posed images, as shown in Fig. 5, and then used to train the 3DGS for detection. Tab. 4.3.1 presents the detection performance for 3DGS trained with the different priors. As reported in Tab. 4.3.1, the 3DGS-DET method using boundary guidance achieved 56.7% in mAP@0.25 and 36.9% in mAP@0.5, demonstrating significant superiority over the center point and mask priors.
@@
4.3.3 Analysis on Different Sampling Methods
In this section, we compare two additional sampling methods with our Box-Focused Sampling: 1) Random Sampling and 2) Farthest Point Sampling (Qi et al., 2017b). The latter iteratively selects points farthest from those already chosen, ensuring even distribution for better scene coverage, focusing on global distribution rather than specific geometric features of objects. The results in Tab. 4.3.1 demonstrate that our Box-Focused Sampling achieves the highest performance, with mAP@0.25 and mAP@0.5 reaching 59.9% and 37.8%, respectively. This is because 3DGS often contain excessive background blobs. Our Box-Focused Sampling is specifically designed to preserve more object-related blobs while suppressing noisy background blobs. In contrast, other sampling methods primarily focus on global scenes without differentiation between objects and background blobs.
@@ 5 Conclusion
In this work, we introduce 3D Gaussian Splatting (3DGS) into 3D Object Detection (3DOD) for the first time. We propose 3DGS-DET, a novel approach that leverages Boundary Guidance and Box-Focused Sampling to enhance 3DGS for 3DOD. Our method effectively addresses the inherent challenges of 3DGS in 3D object detection by improving spatial distribution and reducing background noise. By incorporating 2D Boundary Guidance, we achieve clearer differentiation between objects and background, while Box-Focused Sampling retains more object points and minimizes background noise. Our method demonstrates significant improvements, with gains of +5.6 on mAP@0.25 and +3.7 on mAP@0.5 over the basic pipeline. It also outperforms state-of-the-art NeRF-based methods, achieving +6.6 on mAP@0.25 and +8.1 on mAP@0.5 on the ScanNet dataset, and an impressive +31.5 on mAP@0.25 on the ARKITScenes dataset. These results underscore the effectiveness and superiority of our designs.
Appendix A Appendix
A.1 Performance on NeRF-RPN Setting
Table 5: Performance on the NeRF-RPN setting, which targets class-agnostic box detection. Our method significantly outperforms NeRF-RPN in this setting.
Methods mAP@0.25 mAP@0.5
NeRF-RPN (Hu et al., 2023) 55.5 18.4
3DGS-DET (ours) 75.6  (+20.1) 52.3  (+33.9)
In this section, we adapt our 3DGS-DET to the NeRF-RPN Setting (Hu et al., 2023), which targets class-agnostic box detection. To achieve this, we labeled all the ground-truth boxes with a single ‘object’ category and trained 3DGS-DET accordingly. Additionally, NeRF-RPN uses a different train/validation split compared to the official ScanNet dataset, with its validation set overlapping the official ScanNet training set. To address this, we excluded the overlapping parts between the NeRF-RPN test set and the ScanNet official training set from our training data. We then used the remaining scenes for training, and tested on the same validation set provided by NeRF-RPN. As shown in Tab. 5, 3DGS-DET achieved an mAP@0.25 of 75.6% and an mAP@0.5 of 52.3%, significantly outperforming NeRF-RPN (Hu et al., 2023)’s 55.5% and 18.4%. This demonstrates the significant superiority of our method in the class-agnostic setting.
A.2 Future Work
As the first work to introduce 3DGS into 3DOD, our paper mainly focuses on the primary stage of this pipeline: empowering 3DGS for 3DOD. Diverse experiments demonstrate that our designs can lead to significant improvements. Beyond empowering the 3DGS representation, a subsequent detector specifically designed for 3DGS could hold promise in the future. Besides, exploring joint training of 3DGS and the detector is also an interesting direction. We hope our exploration knowledge, open-source codes and data will inspire further research.
Table 6: Comparison of mAP@0.5 across different methods on ScanNet. The first block presents methods that employ non-view-synthesis representations, including point clouds, RGB-D, and multi-view images. The second block lists methods using view-synthesis representations, such as NeRF-based and our 3DGS-based techniques. Our 3DGS-DET significantly surpasses the NeRF-based NeRF-Det by 8.1 points. Among other representations, 3DGS-DET outperforms all except the point-cloud-based methods FCAF3D and CAGroup3D, which benefit from directly using sensor-captured 3D data, specifically point clouds, as input. Note that some methods did not report mAP@0.5 in previous works, resulting in blank entries for these methods.

### v2

@@ 4.1 Experimental setup
Dataset and Metrics: To thoroughly evaluate the performance of our proposed method in indoor 3D detection tasks, we selected two representative datasets: ScanNet [9] and ARKitScene [1]. ScanNet is a large-scale indoor scene dataset containing over 1,500 real-world 3D scanned scenes, encompassing various complex indoor environments such as residential spaces, offices, and classrooms. The ARKitScene dataset is constructed from RGB-D image sequences, offering detailed geometric inform-ation and precise object annotations. For each scene, a maximum of 600 posed images are extracted. The category settings follow the standard 18 categories for ScanNet and 17 categories for ARKitScene. We use mAP@0.25 and mAP@0.5 as the primary evaluation metrics.
@@
Figure 5: Analysis of guidance from different priors: (a) Center Point Guidance, (b) Mask Guidance, and (c) Boundary Guidance. In (a) and (b), the spatial distribution of Gaussian blobs for objects like the chair, trash bin and sink is incomplete and ambiguous. Gaussian blobs trained with Boundary Guidance exhibit a clearer spatial distribution. For detailed analysis, please refer to Sec. 4.3.
Methods mAP@0.25 mAP@0.5
BP 54.3 34.1
BP+BG 56.7 36.9
BP+BG+BS 59.9 37.8
Table 3: Effect of our designs.
@@
Table 4: Guidance from different priors.
@@
Table 5: Different sampling methods.
4.2 Main Results
Quantitative Results. For the ScanNet dataset, we present the mAP@0.25 and mAP@0.5 performances of various methods in Tab. 1 of the main paper and Tab. 3 of the supplementary, respectively. In both Tab. 1 and Tab. 3 of the supplementary, the methods listed in the first two blocks are non-view-synthesis representation-based 3D detection methods: the first block [33, 37, 46] lists approaches based on point clouds, while the second block focuses on multi-view image methods [45, 41, 38]. The third block consists of view-synthesis representation-based 3DOD methods, including NeRF-Det [52], NeRF-Det++ [16] and our proposed 3DGS-DET. NeRF-Det and NeRF-Det++ are the closest work to ours, leveraging Neural Radiance Fields (NeRF). ‘3DGS-DET’ is our full method, utilizing both Boundary Guidance and Box-Focused Sampling as described in Sec. 3.4. As illustrated in Tab. 1 and Tab. 3 of the supplementary, our method significantly outperforms NeRF-Det++ by +6.0 on mAP@0.25 and +7.8 on mAP@0.5, showcasing the superiority of our approach. Besides, our method also clearly surpasses approaches utilizing multi-view images as input, demonstrating the superiority of our 3DGS reconstruction for detection.
Regarding the ARKitScene dataset, considering NeRF-Det is the closest work to ours, we follow the same setup described in the NeRF-Det [52] supplementary materials: ‘In our experiments, we utilize the subset of the dataset with low-resolution images.’ Similarly, we adopt the same subset of the ARKitScenes dataset. Other methods that report performance on ARKitScene use the full dataset, so our 3DGS-DET is only compared with ImVoxelNet, NeRF-Det and NeRF-Det++ under the same conditions. The results in Tab. 2 demonstrate that 3DGS-DET performs better across most categories, achieving an mAP@0.25 of 58.2, which significantly outperforms NeRF-Det++ by +14.9, highlighting the superiority of our method.
Qualitative results. We provide a qualitative comparison with NeRF-Det in Fig. 1 and Fig. 2 of the supplementary. As shown, our method detects more objects in the scene with greater positional accuracy compared to NeRF-Det [52], demonstrating the superiority of our method.
4.3 Analysis on the Proposed Designs
In this section, we demonstrate the effectiveness of our contributions by first presenting the performance of our proposed basic 3DGS detection pipeline and then incrementally incorporating our additional designs to analyze the resulting performance improvements. We further delve into the method and provide additional analyses to for a more comprehensive understanding of our method.
Our Proposed Basic 3DGS Detection Pipeline. As shown in Tab. 5, ‘BP’ represents our basic pipeline, i.e., our proposed detection pipeline utilizing 3DGS, as described in Sec. 3.2. Benefiting from the advantages of 3DGS as an explicit scene representation, our basic pipeline surpasses NeRF-Det++ (54.3 vs. 53.9), underscoring the significance of introducing 3DGS into indoor 3DOD for the first time.
@@ 4.3 Analysis on the Proposed Designs
Effectiveness of Boundary Guidance. In Tab. 5, ‘BP+BG’ incorporates the proposed Boundary Guidance as detailed in Sec. 3.3. Introducing Boundary Guidance into the basic pipeline results in a significant improvement of 2.4 points (56.7 vs. 54.3), demonstrating the effectiveness of the proposed Boundary Guidance. To further explore the impact of Boundary Guidance on 3DGS representations, we present a visual comparison of the spatial distribution of trained Gaussian blobs in Fig. 3 in the supplementary. As we can see, Gaussian blobs trained with Boundary Guidance demonstrate clearer spatial distribution and more distinct differentiation between objects and the background. We also present rendered images from different views by 3DGS trained with Boundary Guidance in Fig. 4 and Fig. 5 in the supplementary. As can be observed, the category-specific boundaries are clearly rendered and show multi-view stability, indicating that the 3D representation has effectively embedded the priors from Boundary Guidance. All these results clearly verify the effectiveness of the proposed Boundary Guidance for 3DGS-Det.
@@
Effectiveness of Box-Focused Sampling. Furthermore, we introduce Box-Focused Sampling detailed in Sec. 3.4, represented by ‘BP+BG+BS’ in Tab. 5. This addition leads to a further performance boost of 3.2 points (59.9 vs. 56.7), proving the effectiveness of Box-Focused Sampling. The visual comparison of sampled Gaussian blobs is shown in Fig. 6 in the supplementary. We can observe that the proposed Box-Focused Sampling significantly retains more object blobs and suppresses noisy background blobs.
@@
Guidance from Different Priors. In this section, we analyze the impact of guidance from various priors. As described in Sec. 3.3, we utilize the object’s boundary as the guidance prior. Here, we perform an ablation study considering the object’s center point and mask as alternative priors. To obtain the center point, we detect the object’s bounding box using GroundingDINO [22] and compute its center coordinates. The mask is generated with GroundedSAM [36]. Note that all priors are category-specific, with each class associated with a fixed color. These priors are overlaid on the posed images, as shown in Fig. 5, and then used to train the 3DGS for detection. Tab. 5 presents the detection performance for 3DGS trained with the different priors. As reported in Tab. 5, the 3DGS-DET method using boundary guidance achieved 56.7% in mAP@0.25 and 36.9% in mAP@0.5, demonstrating significant superiority over the alternative priors.
Sampling Methods mAP@0.25 mAP@0.5
Box-Focused Sampling 59.9 37.8
Box-Focused Sampling-GT-Box 76.5 68.3
Table 6: Upper bound of Box-Focused Sampling with GT boxes.
Sampling Methods Average Noisy Points
Farthest Point Sampling (FPS) 80025
Box-Focused Sampling (Ours) 59476 (-25.7%)
Table 7: Average noisy points: FPS vs. Box-Focused Sampling.
@@
Different Sampling Methods. In this section, we compare two additional sampling methods with our Box-Focused Sampling: 1) Random Sampling and 2) Farthest Point Sampling [31]. The latter iteratively selects points farthest from those already chosen, ensuring even distribution for better scene coverage, focusing on global distribution rather than specific geometric features of objects. The results in Tab. 5 demonstrate that our Box-Focused Sampling achieves the highest performance, with mAP@0.25 and mAP@0.5 reaching 59.9% and 37.8%, respectively. This is because 3DGS often contain excessive background blobs. Our Box-Focused Sampling is specifically designed to preserve more object-related blobs while suppressing noisy background blobs. In contrast, other sampling methods primarily focus on global scenes without differentiation between objects and background blobs.
The Upper Bound of Box-Focused Sampling. In this section, we conduct an ablation study on the upper bound of Box-Focused Sampling by utilizing ground-truth (GT) 3D bounding boxes to sample 3DGS blobs. As shown in Tab. 6, despite the significant improvements achieved with Box-Focused Sampling (3.2 points in Tab. 5), using GT 3D boxes brings further enhancements, highlighting its potential as a promising research direction.
Class-agnostic Setting mAP@0.25 mAP@0.5
NeRF-RPN [15] 55.5 18.4
NeRF-MAE [17] 57.1 17.0
NeRF-FCM [13] 58.8 23.4
Gaussian-Det [54] 71.7 24.5
3DGS-DET (ours) 75.6  (+3.9) 52.3  (+27.8)
Table 8: Performance on the class-agnostic setting [15, 17, 13, 54], which targets class-agnostic box detection, i.e., localizing objects without predicting their semantic categories.
Noise Reduction via Box-Focused Sampling. We compare the average number of noisy points in test scenes on ScanNet when using Farthest Point Sampling (FPS) and our Box-Focused Sampling. As shown in Tab. 7, our Box-Focused Sampling significantly reduces noisy background blobs by 25.7% compared with the FPS sampling, demonstrating its effectiveness in noise reduction.
4.4 Performance on NeRF-RPN Setting
In this section, we adapt our 3DGS-DET to the NeRF-RPN setting [15, 17], which targets class-agnostic box detection. To achieve this, we labeled all the ground-truth boxes with a single ‘object’ category and trained 3DGS-DET accordingly. Additionally, NeRF-RPN uses a different train/validation split compared to the official ScanNet dataset, with its validation set overlapping the official ScanNet training set. To address this, we excluded the overlapping parts between the NeRF-RPN validation set and the ScanNet official training set from our training data. We then used the remaining scenes for training, and evaluated on the same validation set provided by NeRF-RPN. As shown in Tab. 8, 3DGS-DET achieved an mAP@0.25 of 75.6% and an mAP@0.5 of 52.3%, significantly outperforming Gaussian-Det [54]’s 71.7% and 24.5%. This demonstrates the significant superiority of our method in the class-agnostic setting.
4.5 More Ablation Studies
To comprehensively analyze our method, in Sec. 2 of the supplementary material, we further study ‘The Effect of Boundary Guidance on the Spatial Distribution of 3DGS Blobs’, ‘Comparison with SfM-Based 3D Detection Baseline’, ‘Application to Feed-Forward 3DGS’, ‘Efficiency Analysis’ , ‘Failure Cases’ and provide extensive qualitative analysis.
@@ 5 Conclusion
In this work, we introduce 3DGS into indoor 3D Object Detection (3DOD) for the first time and propose 3DGS-DET, a novel approach that leverages Boundary Guidance and Box-Focused Sampling to enhance 3DGS for indoor 3DOD. Our method addresses 3DGS challenges by improving spatial distribution and reducing background noise. With 2D Boundary Guidance, we achieve clearer object-background differentiation, while Box-Focused Sampling retains more object points and suppresses noise. 3DGS-DET significantly outperforms state-of-the-art methods.
Supplementary Material
1 Implementation Details
To train 3DGS, we follow [18] to initialize the 3D coordinates of Gaussian blobs using Structure-from-Motion (SfM) points. The training hyperparameters are the same as those in [18]. We employ pretrained GroundedSAM [36] and the Suzuki-Abe algorithm [44] as the boundary detector in Boundary Guidance. The pretrained GroundingDINO [22] is used as the box detector in the Box-Focused Sampling strategy. Note that GroundingDINO is integrated as part of GroundedSAM. Therefore, in practice, only GroundedSAM needs to be executed to obtain both predictions. For the detection tool, we utilize FCAF3D [37] implemented in MMDetection3D [8]. The training hyperparameters are the same as those in FCAF3D. In our ablation study, to ensure a fair comparison, all model versions are trained with the same hyperparameters, such as the same number of epochs, specifically 12 epochs. All the ablation experiments (Sec. 4.3 in the main paper) are conducted on ScanNet.
2 Additional Ablation Studies
2.1 The Effect of Boundary Guidance on the Spatial Distribution of 3DGS Blobs
Quantitative Study. To investigate the effect of Boundary Guidance on the spatial distribution of Gaussian blobs, we conducted an ablation study utilizing the first 100 scenes of ScanNet. Specifically, we evaluated the spatial distribution of Gaussian blobs by calculating the Root Mean Square Error (RMSE) metric [61] between the Gaussian blob positions and the ground-truth point clouds. Without Boundary Guidance, the average RMSE was 56.08%, whereas incorporating Boundary Guidance reduced the RMSE to 53.99%, achieving a significant improvement of 2.09 points. This result demonstrates that Boundary Guidance effectively improves the spatial distribution of Gaussian blobs.
Qualitative Study. In addition to the quantitative evaluation, we provide a qualitative analysis of the spatial distribution of Gaussian blobs in Fig. 3. Gaussian blobs trained with Boundary Guidance demonstrate clearer spatial distributions and more distinct differentiation between objects and the background. These qualitative results visually confirm the quantitative findings above, illustrating the effectiveness of Boundary Guidance in improving the spatial distribution of Gaussian blobs.
2.2 Rendered Images from 3DGS Trained with Boundary Guidance
We present rendered images from different views generated by 3DGS trained with Boundary Guidance in Fig. 4 and Fig. 5. As shown, the category-specific boundaries are clearly rendered and exhibit multi-view consistency, demonstrating that the 3D representation effectively incorporates the priors from Boundary Guidance.
2.3 Qualitative Analysis of Box-Focused Sampling for 3DGS Blobs
Fig. 6 visually demonstrates the effect of Box-Focused Sampling. As shown, Box-Focused Sampling enhances the preservation of object blobs while effectively reducing noisy background blobs.
2.4 Comparison with SfM-based 3D Detection Baseline
We implement the baseline combining SfM [39, 40] points with the FCAF3D [37]. As shown in Tab. 2, our method significantly outperforms the SfM-Detector by 11.1 points regarding mAP@0.25 and 8.1 points regarding mAP@0.5. The performance gap primarily stems from SfM reconstructions’ sparse point distributions. In contrast, our 3DGS with Boundary Guidance and Box-Focused Sampling promote superior spatial distribution of dense Gaussian blobs while suppressing background noise, ultimately yielding significant performance gains.
Method mAP@0.25 mAP@0.5
NeRF-Det [52] 53.3 29.7
NeRF-Det++ [16] 53.9 30.0
3DGS-DET ($F$) 58.5 (+4.6) 31.8 (+1.8)
Table 1: When applied to the feed-forward 3DGS, the performance of our method (‘3DGS-DET (FF)’ still remains significantly ahead.
2.5 Application to Feed-forward 3DGS
Vanilla 3DGS [18] optimizes each training scene individually, whereas the detector is trained on the entire training dataset. This fundamental difference in the optimization process prevents end-to-end training. However, to achieve end-to-end training, our method is flexible to be applied not only to Vanilla 3DGS but also to more recent feed-forward 3DGS methods [6], which are trained on entire datasets. We specifically applied our designs to the feed-forward 3DGS in an end-to-end integration manner and still achieved much better results than NeRF-Det++ [16], as shown in Tab. 1.
2.6 Efficiency analysis
When applied to the feed-forward 3DGS, our method (‘3DGS-DET (FF)’ in Tab. 1) completes the entire training phase in 11 hours and the testing phase in 75 seconds on ScanNet, delivering significant performance improvements over NeRF-Det++ with an acceptable computational cost. All timing statistics were measured on two A800 GPUs.
$P_{\text{obj}}$ mAP@0.25 mAP@0.5
SfM-Detector 48.8 29.7
Ours 59.9 (+11.1) 37.8 (+8.1)
Table 2: Comparison with SfM-based 3D Detection Baseline [39, 40, 37].